OpenStudy (anonymous):

average value of y=tan(x) from (-pi/4) to (pi/4) I know how to do it, but I got it wrong. Can someone explain why?

OpenStudy (anonymous):

average of \[y=\tan(x) \] from \[\left[ \frac{ -\pi }{ 4 } ,\frac{ \pi }{ 4 } \right] \]

OpenStudy (anonymous):

does it help to know that the function is odd?

OpenStudy (anonymous):